OpenStudy (anonymous):

Suppose a matrix A, such that A^2=A and A does not equal 0. Does this imply that A is invertible?

OpenStudy (anonymous):

no not necessary

OpenStudy (anonymous):

What matrix would disprove this? Wouldn't A have to be the identity?

OpenStudy (anonymous):

basic rule of invertiblity of matrice says that it should be singular

OpenStudy (anonymous):

Sorry, I'm confused. What would be a matrix that doesn't have an inverse where A^2=A? I can't find one :$

terenzreignz (terenzreignz):

Sorry... didn't read the question thoroughly

OpenStudy (anonymous):

I know, I thought 0 too, but that's not allowed :$

OpenStudy (phi):

let's say A has an inverse A A = A multiply by Ainv Ainv*A*A= Ainv*A A= I It seems A must be I

OpenStudy (phi):

but we can have matrices A*A= A which are not I, so it seems you can not assume this means A is invertible.

terenzreignz (terenzreignz):

On second thought... what about a square matrix with just a 1 in the upper left corner? That's not invertible

jimthompson5910 (jim_thompson5910):

it says in this page http://en.wikipedia.org/wiki/Idempotent_matrix that "With the exception of the identity matrix, an idempotent matrix is singular"

terenzreignz (terenzreignz):

\[\huge A=\left[\begin{matrix}1 & 0 \\ 0 & 0\end{matrix}\right]\] And I think you'll find that \[\huge AA=A\]

jimthompson5910 (jim_thompson5910):

so if A is not the identity matrix, then A is not invertible

OpenStudy (anonymous):

The A=[1,0 0,1] does work. So it isn't necessarily the identity.

OpenStudy (phi):

you can form A using v * v^T / (v^T v) where v is a vector

OpenStudy (anonymous):

10 00 sorry

jimthompson5910 (jim_thompson5910):

10 00 isn't invertible though because it has a determinant of 0

terenzreignz (terenzreignz):

So... that answers the question? :)

OpenStudy (phi):

A A = A does not imply A is invertible

OpenStudy (phi):

It implies if A is invertible, A is I, otherwise A is not invertible

OpenStudy (anonymous):

So, what matrix would disprove this? What matrix has A^2=A but A isn't invertible?

OpenStudy (phi):

v * v^T / (v^T v)

jimthompson5910 (jim_thompson5910):

10 00 is a good counter-example

terenzreignz (terenzreignz):

?? Doesn't A already do that? \[\huge A=\left[\begin{matrix}1 & 0 \\ 0 & 0\end{matrix}\right]\]